3x-y=-10 4x-y=-4 solve. i think i did it wrong

3x-y=-10
4x-y=-4

solve. i think i did it wrong
Answer #1

Given that, 39-y = -10 ceed and 41-4=u4 5) -Y = -47 love =) -Y = -(4+40) =) ya 4840 we put the value of y in 1st equation 3o